OpenStudy (goformit100):

provide me a trick to multiply 6283 x 71 = ??? fastly...

OpenStudy (goformit100):

@UnkleRhaukus

OpenStudy (goformit100):

@Hero @Agent_Sniffles

hero (hero):

I don't really know any tricks,sorry

OpenStudy (goformit100):

@hartnn

OpenStudy (goformit100):

@ganeshie8 @calculusfunctions

hartnn (hartnn):

6283 x 71= 6283 x 70+6283

OpenStudy (goformit100):

IF WE HAVE TO multiply 7527 * 99999 = ? provide me a trick for it....

OpenStudy (anonymous):

just use wolframulator...i dont know any tricks

OpenStudy (anonymous):

have you ever learnt vedic maths:?

OpenStudy (nubeer):

use calculator.

OpenStudy (anonymous):

7527 * 99999 = 7527*10000 - 7527. basically just find a number such that it will simplify the entire calculation.

OpenStudy (calculusfunctions):

Write each factor in expanded form first: (6000 + 200 + 80 + 3)(70 + 1) Now expand as you would polynomial expressions, and simplify.

OpenStudy (lgbasallote):

multiply 6283 by 1....then multip 6283 by 7....

OpenStudy (anonymous):

use a calculator?